# UniTok: A Unified Tokenizer for Visual Generation and Understanding

### v1

@@ 3.4 Unified MLLM
We proceed to develop a unified multimodal model with UniTok. Particularly, we leverage the unified framework introduced in Liquid (Wu et al., 2024b), which models (discrete-valued) vision and language sequences with a universal next-token prediction loss. But instead of learning the visual codebook from scratch, we reuse code embeddings of UniTok by projecting them to the MLLM token space with an MLP projector. Notably, despite UniTok encodes an image into $H\times W\times K$ codes (where $K$ represents the number of sub-codebooks), we simplify this for MLLM input by merging every $K$ consecutive codes into a single visual token. Similarly, when it comes to visual token prediction, we make each token autoregressively predict the next $K$ codes, using a depth transformer head as implemented in RQ-Transformer (Lee et al., 2022) and VILA-U (Wu et al., 2024d). This design maintains efficiency for visual generation in the context of multi-codebooks.
@@ 4.1 Implementation Details
Tokenizer Setup. Leading VQVAE tokenizers predominantly adopt the CNN architecture, while ViT is preferred in CLIP training for its scalability. To take advantage of both, we choose a hybrid architecture, ViTamin-L/16 (Chen et al., 2024), to instantiate UniTok. We configure UniTok with eight sub-codebooks, each containing 4,096 code entries and a latent dimension set to 8-d (the global latent dimension is thus 64-d). The discriminator is initialized with pretrained DINOv2-S (Oquab et al., 2023). We train the tokenizer for one epoch on the public dataset DataComp-1B (Gadre et al., 2024) consisting of 1.28B image-text pairs, with all images resized to $256\times 256$ resolution and a global batch size of 16k. The learning rate is set to 1e-3 for the tokenizer and 2e-4 for the discriminator. Besides, we prepare two settings for evaluation: one with pretrained CLIP weight initialization and one with random initialization (the default setting).
MLLM Setup. We instantiate a unified MLLM described in Section 3.4 with the Llama-2-7B base model (Touvron et al., 2023). Following Liquid, we first pretrain the model on a mix of multimodal data, which is composed of 10M language data from DCLM (Li et al., 2024b), 30M internal MidJourney-style synthetic data, and 30M re-captioned image-text pairs from COYO (Minwoo et al., 2022) and Laion (Schuhmann et al., 2022). Subsequently, we finetune the model on 1.5M text-to-image data and 1.5M multimodal instruction tuning data introduced in Mini-Gemini (Li et al., 2024d). Specifically, the learning rate is set to 5e-5 in the pretraining stage and 2e-5 in the finetuning stage. For visual understanding evaluation, we report results on standard VQA benchmarks including VQAv2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), TextVQA (Singh et al., 2019), POPE (Li et al., 2023), MME (Yin et al., 2023), and MM-Vet (Yu et al., 2023d). For visual generation evaluation, we report results on GenAI-Bench (Lin et al., 2025) and MJHQ-30K (Li et al., 2024a).
@@ 4.2 Tokenizer Comparison
Table 1: Comparison with modern tokenizers on ImageNet reconstruction FID and zero-shot classification accuracy. The rFID is measured at $256\times 256$ resolution with 16×16$\times$16 × downsample ratio. $\dagger$ indicates the model uses pretrained CLIP weights for initialization.
@@
CLIP (Radford et al., 2021) 256 76.2
SigLIP (Zhai et al., 2023) 256 80.5
ViTamin (Chen et al., 2024) 256 81.2
@@
TokenFlow (Qu et al., 2024) 680 1.37
VILA-U (Wu et al., 2024d) 256 1.80 73.3
UniTok 256 0.39 70.5
@@
Emu (Sun et al., 2023) Llama-13B Continuous 224 52.0 - - - - -
LaVIT (Jin et al., 2023) Llama-7B Continuous 224 66.0 46.8 - - - -
DreamLLM (Dong et al., 2023) Vicuna-7B Continuous 224 72.9 - 41.8 - - 26.6
Unified-IO 2 (Lu et al., 2024) 6.8B from scratch Continuous 384 79.4 - - 87.7 - -
Janus (Wu et al., 2024a) DeepSeek-1.3B Continuous 384 77.3 59.1 - 87.0 1338 34.3
CM3Leon (Yu et al., 2023c) 7B from scratch Discrete 256 47.6 - - - - -
LWM (Liu et al., 2024b) Llama-2-7B Discrete 256 55.8 44.8 18.8 75.2 - -
Show-o (Xie et al., 2024) Phi-1.5-1.3B Discrete 256 59.3 48.7 - 73.8 948 -
Chameleon (Team, 2024) 34B from scratch Discrete 512 69.6 - - - -
Liquid (Wu et al., 2024b) Gemma-7B Discrete 512 71.3 58.4 42.4 81.1 1119 -
VILA-U (Wu et al., 2024d) Llama-2-7B Discrete 256 75.3 58.3 48.3 83.9 1336 27.7
@@
SD v2.1 (Rombach et al., 2022a) Diffusion 2000M 0.68 0.70 0.68 0.54 0.64 0.62
SD-XL (Podell et al., 2023) Diffusion 2000M 0.71 0.73 0.69 0.50 0.66 0.63
Midjourney v6 (Radhakrishnan, 2023) Diffusion 0.78 0.78 0.79 0.50 0.76 0.69
DALL-E 3 (Betker et al., 2023) Diffusion 0.82 0.78 0.82 0.48 0.80 0.70
Show-o (Xie et al., 2024) Discrete Diff. 36M 0.70 0.62 0.71 0.51 0.65 0.60
LWM (Liu et al., 2024b) Autoregressive 0.59 0.58 0.54 0.49 0.52 0.53
VILA-U (Wu et al., 2024d) Autoregressive 15M 0.70 0.71 0.74 0.53 0.66 0.64
Liquid (Wu et al., 2024b) Autoregressive 30M 0.76 0.73 0.74 0.46 0.74 0.65
@@
We benchmark UniTok on ImageNet using two primary metrics: Fréchet Inception Distance (FID) to evaluate reconstruction quality, and top-1 zero-shot accuracy to assess image-text alignment. The results are presented in Table 1. Notably, UniTok excels in reconstruction quality compared to both unified and domain-specific tokenizers, recording an impressive 0.38 rFID on ImageNet with 16×16$\times$16 × downsampling ratio. As a discrete tokenizer, UniTok even surpasses the continuous VAE tokenizer from Stable Diffusion v2.1 (Rombach et al., 2022b), showcasing the superiority of the proposed multi-codebook quantization. For the perception performance, we observe that randomly initialized UniTok demonstrates suboptimal zero-shot classification accuracy. This is expected as current training schedule (i.e., one epoch on 1.28B samples) is insufficient for CLIP training to fully converge. It can be seen that initializing the model with pretrained CLIP weights largely alleviates the problem, boosting the zero-shot accuracy from 70.8% to 78.6%. However, we would like to point out that higher ImageNet accuracy does not guarantee superior downstream performance. As demonstrated in Table 7, we find that random initialization actually leads to better understanding performance.
In complement to quantitative results, we also provide examples of reconstructed images in Figure 6.
4.3 Visual Understanding Performance
We evaluate the understanding performance of UniTok on diverse VQA benchmarks in Table 2. Our unified MLLM showcases clear advantages when compared to other unified models that also utilize a discrete visual tokenizer. Specifically, UniTok significantly outperforms the Chameleon model, which relies on a traditional VQVAE tokenizer, by 7.2% higher accuracy on VQAv2. Additionally, it surpasses VILA-U, another model with a unified tokenizer, by 3.3% in accuracy on the TextVQA benchmark and by a notable margin of 112 points on the MME-Perception scores. Furthermore, we can see that UniTok largely narrows the performance gap with MLLMs that incorporate continuous visual tokenizers. These strong results confirm the candidacy of UniTok as a unified visual tokenizer for multimodal models.
4.4 Visual Generation Performance
@@
SD-XL (Podell et al., 2023) Diffusion 1024 9.55
PixArt (Chen et al., 2023) Diffusion 1024 6.14
Playground (Li et al., 2024a) Diffusion 1024 4.48
Liquid (Wu et al., 2024b) Autoregressive 512 5.47
Janus (Wu et al., 2024a) Autoregressive 384 10.10
LWM (Liu et al., 2024b) Autoregressive 256 17.77
Show-o (Xie et al., 2024) Discrete Diff. 256 15.18
VILA-U (Wu et al., 2024d) Autoregressive 256 12.81
@@
Quantitative Results. Table 3 presents the text-to-image generation performance of our unified MLLM on GenAI-Bench (advanced prompts) [^1] Due to paper length limit, we put the results on GenAI-Bench (basic prompts) in the Appendix., a challenging benchmark that measures the alignment of generated images with text prompts in complex dimensions such as counting, differentiation, comparison, and understanding logical relationships. In this demanding context, our model not only consistently outperforms other autoregressive unified models, but also achieves competitive performance against domain experts (diffusion models) trained on billions of images. The strong results underscore the superior capability of our unified MLLM in complex text-to-image generation tasks.
We further evaluate the quality of images generated by our model on the MJHQ-30K benchmark, details of which are presented in Table 4. Notably, as this benchmark primarily relies on the FID score for evaluation, high-resolution images are preferred because they potentially capture more fine-grained details. Despite this makes FID across different resolutions less comparable, we show that our model achieves impressive performance even at the the smallest resolution, showcasing its ability to generate high-quality, detail-rich images.
Qualitative Results. We present some examples of the images generated by our model in Figure 5, using text prompts sampled from MJHQ-30K. The visualization results demonstrate our model is capable of synthesizing photo-realistic and visually appealing images. Moreover, the model is able to comprehend a wide spectrum of concepts, such as ‘Vincent van Gogh painting style’ and ‘bitcoin’, and flexibly combine these concepts to synthesize creative images.
@@ 4.5 Ablation Studies
Table 5: Impact of different supervision types on downstream generation and understanding performance. The rFID and gFID are measured on the ImageNet ($256\times 256$) validation set. LlamaGen-L (Sun et al., 2024a) is adopted as the generator for gFID evaluation.
@@
Number of Sub-Codebooks. To gain deeper insights into multi-codebook quantization, we evaluate how tokenizer performance changes with the number of sub-codebooks. Specifically, for rFID evaluation, we train the tokenizer solely with reconstruction loss on OpenImages (Kuznetsova et al., 2020), and evaluated it on ImageNet ($256\times 256$) validation set. While for ImageNet zero-shot accuracy evaluation, the tokenizer is trained on DataComp-1B 128m subset using only contrastive loss. As shown in Table 6, given a constant global codebook size, increasing the number of sub-codebooks consistently improves reconstruction FID and classification accuracy. This indicates that multi-codebook quantization generally benefits vector-quantized models, independent of the training objectives.
Table 6: Ablation on the number of sub-codebooks. The size of a codebook is denoted as $A\times B$, where $A$ is the number of sub-codebook and $B$ is the size of sub-codebook.

### v2

@@ 3.4 Unified MLLM
We proceed to develop a unified multimodal model with UniTok. Particularly, we leverage the unified framework introduced in Liquid [60], which models (discrete-valued) vision and language sequences with a universal next-token prediction loss. But instead of learning the visual codebook from scratch, we reuse code embeddings of UniTok by projecting them to the MLLM token space with an MLP projector. Notably, despite UniTok encodes an image into $H\times W\times K$ codes (where $K$ represents the number of sub-codebooks), we simplify this for MLLM input by merging every $K$ consecutive codes into a single visual token. Similarly, when it comes to visual token prediction, we make each token autoregressively predict the next $K$ codes, using a depth transformer head as implemented in RQ-Transformer [18] and VILA-U [62]. This design maintains efficiency for visual generation in the context of multi-codebooks.
@@ 4.1 Implementation Details
Tokenizer Setup. Leading VQVAE tokenizers predominantly adopt the CNN architecture, while ViT is preferred in CLIP training for its scalability. To take advantage of both, we choose a hybrid architecture, ViTamin-L/16 [5], to instantiate UniTok. We configure UniTok with eight sub-codebooks, each containing 4,096 code entries and a latent dimension set to 8-d (the global latent dimension is thus 64-d). The discriminator is initialized with pretrained DINOv2-S [35]. We train the tokenizer for one epoch on the public dataset DataComp-1B [9] consisting of 1.28B image-text pairs, with all images resized to $256\times 256$ resolution and a global batch size of 16k. The learning rate is set to 1e-3 for the tokenizer and 2e-4 for the discriminator. Besides, we prepare two settings for evaluation: one with pretrained CLIP weight initialization and one with random initialization (the default setting).
MLLM Setup. We instantiate a unified MLLM described in Section 3.4 with the Llama-2-7B base model [53]. Following Liquid, we first pretrain the model on a mix of multimodal data, which is composed of 10M language data from DCLM [22], 30M internal MidJourney-style synthetic data, and 30M re-captioned image-text pairs from COYO [33] and Laion [42]. Subsequently, we finetune the model on 1.5M text-to-image data and 1.5M multimodal instruction tuning data introduced in Mini-Gemini [23]. Specifically, the learning rate is set to 5e-5 in the pretraining stage and 2e-5 in the finetuning stage. For visual understanding evaluation, we report results on standard VQA benchmarks including VQAv2 [13], GQA [14], TextVQA [43], POPE [24], MME [65], and MM-Vet [71]. For visual generation evaluation, we report results on GenAI-Bench [26] and MJHQ-30K [20].
@@ 4.2 Tokenizer Comparison
Table 1: Comparison on ImageNet reconstruction FID and zero-shot classification accuracy. rFID is measured at 256$\times$256 resolution with 16×16$\times$16 × downsample ratio. $\dagger$ indicates model using pretrained CLIP weights for initialization. $\ast$ indicates model trained on OpenImages.
@@
CLIP [39] 256 76.2
SigLIP [72] 256 80.5
ViTamin [5] 256 81.2
@@
TokenFlow [38] 680 1.37
VILA-U [62] 256 1.80 73.3
UniTok 256 0.41 70.8
@@
We benchmark UniTok on ImageNet using two primary metrics: Fréchet Inception Distance (FID) to evaluate reconstruction quality, and top-1 zero-shot accuracy to assess image-text alignment. The results are presented in Table 1. To provide a fair comparison with tokenizers trained on small datasets, we also train a version of UniTok on OpenImages [17] solely with reconstruction supervision. It can be seen that UniTok excels in reconstruction quality compared to both unified and domain-specific tokenizers, recording an impressive 0.38 rFID on ImageNet with 16×16$\times$16 × downsampling ratio. As a discrete tokenizer, UniTok even surpasses the continuous VAE tokenizer from Stable Diffusion v2.1 [41], showcasing the superiority of the proposed multi-codebook quantization. For the perception performance, we observe that randomly initialized UniTok demonstrates suboptimal zero-shot classification accuracy. This is expected as current training schedule (i.e., one epoch on 1.28B samples) is insufficient for CLIP training to fully converge. It can be seen that initializing the model with pretrained CLIP weights largely alleviates the problem, boosting the zero-shot accuracy from 70.8% to 78.6%. In complement to quantitative results, we provide examples of reconstructed images in Figure 6.
4.3 Class-Conditional Image Generation
Recent studies on diffusion models indicate that injecting semantics into VAE training leads to a better-structured latent space, considerably enhancing guidance-free generation performance [64, 4, 19]. To evaluate whether UniTok possesses similar properties, we test it within the LlamaGen framework for class-conditional image generation. As shown in Table 2, UniTok reduces the FID by 12.11 compared to the VQGAN baseline in CFG-free generation, under the same generator setup. This implies that UniTok learns a more structured code distribution, benefiting autoregressive modeling.
Table 2: Class-conditional image generation results on ImageNet 256$\times$256. : VQGAN tokenizer from LlamaGen. : Images are generated at 384$\times$384 resolution and then resized to 256$\times$256 for evaluation. ‘Pre.’: precision; ‘Rec.’: recall; ‘CFG’: classifier-free-guidance.
Tokenizer rFID Generator #Params. Generation w/o CFG Generation w/ CFG
[…]
Diffusion Models
SD-VAE [41] 0.61 DiT [36] 675M 9.62 121.5 0.67 0.67 2.27 278.2 0.83 0.57
SiT [31] 675M 8.61 131.7 0.68 0.67 2.06 270.3 0.82 0.59
Masked Generative Models
LFQ [68] 0.9 MAGVIT-v2 [68] 307M 3.07 213.1 1.91 324.3
TiTok-L [70] 2.21 MaskGIT [3] 177M 3.15 173.0 2.77 199.8
Autoregressive Models
VQGAN 2.19 LlamaGen [44] 1.4B 14.65 86.3 0.63 0.68 2.34 253.9 0.81 0.60
UniTok (Ours) 0.41 LlamaGen [44] 1.4B 2.51 216.7 0.82 0.57 2.77 227.5 0.81 0.57
4.4 Unified Understanding and Generation
Understanding Performance. We evaluate the understanding performance of UniTok on diverse VQA benchmarks in Table 3. Our unified MLLM showcases clear advantages when compared to other unified models that also utilize a discrete visual tokenizer. Specifically, UniTok significantly outperforms the Chameleon model, which relies on a traditional VQVAE tokenizer, by 7.2% higher accuracy on VQAv2. Additionally, it surpasses VILA-U, another model with a unified tokenizer, by 3.3% in accuracy on the TextVQA benchmark and by a notable margin of 112 points on the MME-Perception scores. Furthermore, we can see that UniTok largely narrows the performance gap with MLLMs that incorporate continuous visual tokenizers. These strong results confirm the candidacy of UniTok as a unified visual tokenizer for multimodal models.
@@ 4.4 Unified Understanding and Generation
Emu [46] Llama-13B Continuous 224 52.0 - - - - -
LaVIT [15] Llama-7B Continuous 224 66.0 46.8 - - - -
DreamLLM [7] Vicuna-7B Continuous 224 72.9 - 41.8 - - 26.6
Unified-IO 2 [29] 6.8B from scratch Continuous 384 79.4 - - 87.7 - -
Janus [59] DeepSeek-1.3B Continuous 384 77.3 59.1 - 87.0 1338 34.3
CM3Leon [69] 7B from scratch Discrete 256 47.6 - - - - -
LWM [27] Llama-2-7B Discrete 256 55.8 44.8 18.8 75.2 - -
Show-o [63] Phi-1.5-1.3B Discrete 256 59.3 48.7 - 73.8 948 -
Chameleon [47] 34B from scratch Discrete 512 69.6 - - - -
Liquid [60] Gemma-7B Discrete 512 71.3 58.4 42.4 81.1 1119 -
VILA-U [62] Llama-2-7B Discrete 256 75.3 58.3 48.3 83.9 1336 27.7
@@
Generation Performance. Table 4 presents the text-to-image generation performance of our unified MLLM on GenAI-Bench (advanced prompts), a challenging benchmark that measures the alignment of generated images with text prompts in complex dimensions such as counting, differentiation, comparison, and understanding logical relationships. In this demanding context, our model not only consistently outperforms other autoregressive unified models, but also achieves competitive performance against domain experts (diffusion models) trained on billions of images. The strong results underscore the superiority of our unified MLLM in complex text-to-image generation tasks. We also provide results on GenAI-Bench (basic prompts) in Table 9 in Appendix.
@@
SD v2.1 [40] Diffusion 2000M 0.68 0.70 0.68 0.54 0.64 0.62
SD-XL [37] Diffusion 2000M 0.71 0.73 0.69 0.50 0.66 0.63
Midjourney v6 Diffusion 0.78 0.78 0.79 0.50 0.76 0.69
DALL-E 3 [2] Diffusion 0.82 0.78 0.82 0.48 0.80 0.70
Show-o [63] Discrete Diff. 36M 0.70 0.62 0.71 0.51 0.65 0.60
LWM [27] Autoregressive 0.59 0.58 0.54 0.49 0.52 0.53
VILA-U [62] Autoregressive 15M 0.70 0.71 0.74 0.53 0.66 0.64
Liquid [60] Autoregressive 30M 0.76 0.73 0.74 0.46 0.74 0.65
@@
We further evaluate the quality of images generated by our model on the MJHQ-30K benchmark, details of which are presented in Table 5. Notably, as this benchmark primarily relies on the FID score for evaluation, high-resolution images are preferred because they potentially capture more fine-grained details. Despite this makes FID across different resolutions less comparable, we show
@@
SD-XL [37] Diffusion 1024 9.55
PixArt [6] Diffusion 1024 6.14
Playground [20] Diffusion 1024 4.48
Liquid [60] Autoregressive 512 5.47
Janus [59] Autoregressive 384 10.10
LWM [27] Autoregressive 256 17.77
Show-o [63] Discrete Diff. 256 15.18
VILA-U [62] Autoregressive 256 12.81
@@
that our model achieves impressive performance even at the the smallest resolution, showcasing its ability to generate high-quality, detail-rich images.
We present some examples of the images generated by our model in Figure 4, using text prompts sampled from MJHQ-30K. The visualization results demonstrate our model is capable of synthesizing photo-realistic and visually appealing images. Moreover, the model is able to comprehend a wide spectrum of concepts, such as ‘Vincent van Gogh painting style’ and ‘bitcoin’, and flexibly combine these concepts to synthesize creative images.
@@ 4.5 Ablation Studies
Table 6: Impact of different supervision types on downstream generation and understanding performance. The rFID and gFID are measured on the ImageNet ($256\times 256$) validation set. LlamaGen-L [44] is adopted as the generator for gFID evaluation.
@@
MCQ v.s. RQ. Following discussions in Section 3.3, we provide an apple-to-apple comparison between multi-codebook quantization and residual quantization in Table 8. For fair comparisons, we directly implement RQ on our codebase, keeping all the training settings the same as UniTok. Both tokenizers are trained on a 512M subset of DataComp-1B. Notably, unlike MCQ, RQ by default uses a shared large codebook. To keep the global codebook size the same as UniTok, we set the codebook size of RQ to 32768. It can be seen that RQ demonstrates inferior reconstruction performance and lower classification accuracy compared to MCQ under the high bottleneck dimension (64-d) setting.
Number of Sub-Codebooks. To gain deeper insights into multi-codebook quantization, we evaluate how tokenizer performance changes with the number of sub-codebooks in Table 8. Specifically, the size of a codebook is denoted as $A\times B$, where $A$ is the number of sub-codebook and $B$ is the size of sub-codebook. For rFID evaluation, we train the tokenizer solely with reconstruction loss on OpenImages [17], and evaluated it on ImageNet ($256\times 256$) validation set. While for ImageNet zero-shot accuracy evaluation, the tokenizer is trained on DataComp-1B 128m subset using only contrastive loss. Given a constant global codebook size, we see that increasing the number of sub-codebooks consistently improves reconstruction FID and classification accuracy. This indicates that multi-codebook quantization generally benefits vector-quantized models, independent of the training objectives.
Table 7: Comparison between MCQ and RQ.
[…]
RQ 1$\times$32768 64 3.46 58.8
MCQ 8$\times$4096 64 0.55 63.7
Table 8: Ablation on sub-codebook number.
@@
CLIP Weight Initialization. We notice that higher ImageNet accuracy does not guarantee superior downstream performance. In the setting of UniTok, random initialization actually leads to better understanding performance. Due to paper length limit, we put discussion of this part in the Appendix.
5 Limitations and Conclusion
This paper studies unified visual tokenization for generation and understanding, which serves as the cornerstone of unified multimodal large language models. We investigate the training paradigm of unified tokenizers and identify that the current challenge in unification mainly arises from the limited representational power of discrete tokens. To address this limitation, we introduce multi-codebook quantization and attention projection to build a unified tokenizer called UniTok. We show that UniTok excels in downstream visual generation and understanding tasks. The ablation study further reveals that discriminative and generative representation learning does not inherently conflict. We hope our findings could inspire future research in this domain.
Appendix A Attention Projection Modules
Figure 5 illustrates the adaptations we made to the tradition MHA module, which enables channel compression and expansion in token factorization.
Figure 5: Modified attention blocks for factorization. Modules in yellow indicate a change in the number of channels. $C$ and $c$ stand for the channel dimension, $h$ is the number of heads in the multi-head attention module. $C=h\times c$.
Appendix B Image Reconstruction
Appendix C More Generation Results
[…]
Spatial Action Part
SD v2.1 [40] Diffusion 2000M 0.80 0.79 0.76 0.77 0.80 0.78
SD-XL [37] Diffusion 2000M 0.84 0.84 0.82 0.83 0.89 0.83
Midjourney v6 [midjourney] Diffusion 0.88 0.87 0.87 0.87 0.91 0.87
DALL-E 3 [2] Diffusion 0.91 0.90 0.92 0.89 0.91 0.90
Show-o [63] Discrete Diff. 36M 0.72 0.72 0.70 0.70 0.75 0.70
LWM [27] Autoregressive 0.63 0.62 0.65 0.63 0.70 0.63
VILA-U [62] Autoregressive 15M 0.78 0.78 0.77 0.78 0.79 0.76
Liquid [60] Autoregressive 30M 0.84 0.86 0.81 0.83 0.91 0.83
UniTok Autoregressive 30M 0.85 0.87 0.86 0.86 0.89 0.85
Appendix D CLIP Weight Initialization
In Table 10, we ablate the impact of CLIP weight initialization on visual understanding performance. Specifically, we adopt the classic LLaVA framework for evaluation, replacing the original CLIP tokenizer with UniTok while keeping all other the training settings unchanged. One tokenizer is initialized with the pretrained ViTamin-L-256 [5] weights, while the other is randomly initialized. To our surprise, UniTok that is trained from scratch surpasses the one initialized with pretrained CLIP weights, despite the latter actually achieves better zero-shot classification accuracy. This suggests downstream VQA performance may not be highly correlated with ImageNet classification accuracy. More importantly, it also implies that CLIP weight initialization may serve as a negative prior for unified tokenizers, as the unified visual feature space could drastically differ from CLIP feature space.